1. The present ages of Ann and Annie are 15 and 24 years. x years ago, Annie was twice as old as Ann was. How many years ago was that?

Thank you very much for the answer ! :)

Answer by edjones(8007)

(24-x)/(15-x)=2
24-x=2(15-x)
24-x=30-2x
x=6
.
Ed